OpenStudy (anonymous):

|Medal|Fan| Which example supports the given conjecture? The sum of three odd numbers is an odd number. A. 3 + 5 + 7 = 15 B. 1 + 2 + 3 = 6 C. 2 + 4 + 8 = 14 D. 4 + 5 + 6 = 15

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Help :(

OpenStudy (anonymous):

3,5,7 are odd numbers. 15 is a odd number. 3+5+7=15 Your answer should be A. @ShootingStarz

OpenStudy (anonymous):

Thanks Can ya help me with 2 more please?

OpenStudy (anonymous):

What are the questions.

OpenStudy (anonymous):

Which number supports the given conjecture? If the sum of the digits of a number is divisible by 9, then the number is also divisible by 9. A. 19 B. 54 C. 87 D. 109

OpenStudy (anonymous):

54/9=6 6*9=54 B. is your answer. since the sum of the digit is divisuble by 9, and the number is also divisible by 9.

OpenStudy (anonymous):

The other numbers are all decimals too so that would be your only choice when they are divided by 9.

OpenStudy (anonymous):

Thanks so much i just need help with one more so if you could help that'd be great! :)

OpenStudy (anonymous):

Which equation supports the given conjecture? The sum of the squares of two consecutive integers is an odd integer. A. (2 + 3)2 B. 32 + 52 = 34 C. 4 + 52 = 29 D. 52 + 62 = 61

OpenStudy (anonymous):

@YoungStudier can you help :D

OpenStudy (anonymous):

It's so hard T_T

OpenStudy (anonymous):

I've seen worst then this. I'll take a crack at it.

OpenStudy (anonymous):

lol, But i have no idea what to put.

OpenStudy (anonymous):

Ok The options are actually \[A.(2+3)^2\] \[B.3^2+5^2=34\] \[C.4+5^2=29\] \[D.5^2+6^2=61\] @ShootingStarz Please put the right answers out please next time.

OpenStudy (anonymous):

It's okay the right answer is D And the answers are correct :)

OpenStudy (anonymous):

I already figured out what it was. :)

OpenStudy (anonymous):

Please do not just copy and paste. Try to make the answers like this when they show up. If you copy and paste they don't even look right. @ShootingStarz

OpenStudy (anonymous):

Oh, okay Sorry.

OpenStudy (anonymous):

You need anymore help with other questions? @ShootingStarz

OpenStudy (anonymous):

No thanks i'm done i got a 100% :D

OpenStudy (anonymous):

Thanks for asking.

OpenStudy (anonymous):

Good job. @ShootingStarz Your getting good at math I see?

OpenStudy (anonymous):

Thanks, and yeah i guess..i just did some research lol :D

OpenStudy (anonymous):

Thanks for helping me, and i gtg Byeee! :D